Shift Summary for Owl Shift 8/16/99

S. Elizabeth Turner (eturner@kudzu)
Mon, 16 Aug 1999 07:01:59 -0500

Shift Summary for owl shift, 16/AUG/99

On Shift: SET, JL

Hours of beam available at all intensities: 6
Hours of beam available at nominal intensity: 6
Hours of E832 data (special runs not included): 6
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time:
TeV ramp trip (~45 min), energy conservation (~15 min)
========================================================================
Runs ended this shift:
none

run 14236 (BEAM_832) still running at end of shift

Tapes brought to FCC this shift: KQ1001, ...
(including tapes for online split)
none

========================================================================
Accesses during shift:
None.
========================================================================
Overview of shift:
E.g., calibrate CsI, E832 datataking

Very smooth 832 datataking.
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14236 162 2:20 lose beam -- TeV ramp trip
194 3:05 beam returns
203 ktevana errors -- kfrend
207 ktevana errors -- kfrend
4:28 Notice that an L3 driver process has died
on ktev3, doesn't appear to affect data-taking.
292 5:15 lose beam -- energy conservation
297 beam returns
307 lose beam
314 beam returns
========================================================================
Follow up Notes on Previous Shift
none
========================================================================
Items to Follow up on Next Shift

Do drt_fresh_start and fix_disp at tape change (to restart L3 driver
process and fix DC histograms which are now being run by fix_disp).